# DeepInteraction++: Multi-Modality Interaction for Autonomous Driving

### v2

@@ I Introduction
To overcome the aforementioned limitations, in this work a novel modality interaction strategy, termed DeepInteraction++, for integrating information from different sensors is introduced (Figure 1(b)). Our key idea is to learn and maintain multiple modality-specific representations instead of deriving a single fused representation. This approach enables inter-modality interaction, allowing for the spontaneous exchange of information and the retention of modality-specific strengths with minimal interference between them. Specifically, we start by mapping 3D point clouds and 2D multi-view images into the multi-scale LiDAR BEV features and camera panoramic features using two separate feature backbones in parallel. Subsequently, with an encoder we interact the heterogeneous features for progressive representation learning and integration in a bilateral manner. To fully exploit per-modality representations, we design a decoder to conduct multi-modal predictive interaction in a cascaded manner to yield more accurate perception results.
@@ A.2 Multi-modal representational interaction (MMRI)
where $\boldsymbol{h}^{[i,j]}$ denotes indexing the element at location $(i,j)$ on the 2D representation $\boldsymbol{h}$, and $f_{{\phi}_{p\rightarrow c}}\left(\boldsymbol{h}_{c},\boldsymbol{h}_{p}\right)$ is LiDAR-to-image representational interaction (MMRI-I2L), which produces the image feature map augmented with the LiDAR information.
@@
Furthermore, we employ the flash attention [52, 53] to minimize the additional computation and memory overhead introduced by this module. The experimental results in Section V demonstrate that this operation provides a beneficial dense context, which complements the original object-centric sparse interaction. This finally significantly enhances detection performance and enables the extension of our approach to end-to-end planning.
@@ A.3 Intra-modal representational learning (IML)
Beyond directly incorporating information from heterogeneous modalities, intra-modal reasoning is helpful for more comprehensive integration of these representations. Therefore, in each layer of the encoder, we conduct intra-modal representational learning complementary to multi-modal interaction. In this work, we utilize deformable attention [54] for intra-modal representational learning, replacing the stand-alone attention [55] employed in the original DeepInteraction. Considering the scale variance introduced by perspective projection, interaction operation with a more flexible receptive field would be more reasonable than conducting cross-attention within fixed local neighbors as [10]. This modification maintains the original efficient local computation while achieving a more flexible receptive field and facilitating the multi-scale interaction.
@@ A.4 Efficient interaction with grouped sparse attention
To tackle this issue, we carefully examine the distribution of the number of valid image tokens per pillar and divide these pillars into several intervals =$\mathcal{I}=\left\{(N_{i},N_{i+1})\right\}_{i=0}^{N_{\mathrm{interval}}}$. Then we batchify pillars within each interval by padding the number of keys and values to the interval’s upper limit $N_{i+1}$ for attention computations. With careful selection of interval boundaries, this modification significantly reduces memory consumption with negligible impact on parallelism. Additionally, it is computationally equivalent to the original implementation, as the padded tokens are masked during the attention process.
@@ III-B Decoder: Multi-modal predictive interaction
TABLE I: Comparison with state-of-the-art methods for 3D object detection on the nuScenes test set. Metrics: mAP(%), NDS(%).
@@ IV DeepInteraction++ for end2end autonomous driving
To further demonstrate the scalability and superiority, we extend our DeepInteraction++ to an end-to-end multi-task framework, simultaneously resolving scene perception, motion prediction, and ego-planning tasks. Instead of involving several sub-tasks for the comprehensive analysis of the driving scenario, we affiliate three additional downstream tasks (including map segmentation, prediction, and planning) following VAD [47], a relatively lightweight framework. Hence, our end-to-end variant can effectively alleviate the memory overhead due to the complicated interaction encoder and further unleash the multi-task capabilities benefiting from multi-modal representations.
@@ V-A Experimental setup
TABLE II: Quantitative comparison of detection performance between DeepInteraction and DeepInteraction++ under different image backbones. The results are evaluated on the nuScenes val split.
[…]
backbone
DeepInteraction R50 72.63 69.85 26.70 24.95 27.57 24.77 18.94
DeepInteraction++ 72.91 70.12 26.52 25.08 26.62 24.45 18.87
DeepInteraction Swin-Tiny 72.62 69.97 26.96 25.15 28.10 24.46 18.90
DeepInteraction++ 73.28 70.63 26.79 25.25 26.24 23.43 18.61
TABLE III: Run time comparison measured on an NVIDIA RTX A6000 GPU. If not specified with , the performance is evaluated on the nuScenes val set.
@@ V-C Comparison to the state of the arts
Main results. We compare with state-of-the-art alternatives on both the val and test splits of nuScenes dataset. As shown in Table I, our vanilla DeepInteraction has surpassed all its prior arts under the same settings by a considerable margin, and our DeepInteraction++ achieves new state-of-the-art performance with the improved architectural design. Notably, compared to Transfusion [32], which is a representative uniliteral fusion baseline, our DeepInteraction provides a significant performance gain of 2.4% mAP and 1.3% NDS using the same modality-specific backbone and training recipe, verifying the advantages of our multi-modal interaction approach. We provide the per-category results in Table IX. The qualitative results are shown in Figure 4.
Our DeepInteraction++ by default employs a stronger image backbone. To demonstrate that the improvements brought by the revised architecture are consistent and essential, we additionally provide a systematic and comprehensive comparison between DeepInteraction and DeepInteraction++ under the same image backbone on the nuScenes val set. The results in Table II suggest that DeepInteraction++ with a more meticulously designed architecture consistently beats the baseline across most metrics under all settings while adhering to the same hierarchical modality interaction build.
@@
Runtime. We compare the inference speed of all methods on NVIDIA RTX A6000 GPU. As shown in Table III, our method achieves the best performance with faster inference speed than alternative painting-based [2] and query-based [38, 59] fusion approaches. This demonstrates that our method achieves a better trade-off between performance and efficiency. Specifically, feature extraction for multi-view high-resolution camera images contributes the most of the overall latency in a multi-modal 3D detector as verified in [2]. Our interaction modules are built with a relatively lighter model architecture that offers better running speed. From Figure 6, we observe that increasing the number of decoder layers only brings negligible extra latency, which concurs with the same conclusion.
@@ V-D1 Ablations of the modality interaction
Effects of the representational interaction. To demonstrate the superiority of our multi-modal representational interaction, we compare it with a degraded baseline, which does not iteratively refine the image features during the representational interactions. For a fair comparison, both methods use the same number of encoder layers as well as the same decoder. As shown in Table IV a) and c), our representational interaction is more effective than the unilateral fusion alternatives. Additionally, we compare the representative baseline Transfusion [32] with the conventional modality fusion strategy in Tables VIII, indicating the advantages of our bilateral modality interaction strategy.
@@ V-D2 Ablations on the encoder
TABLE V: Ablation on the encoder design. IML: Intra-modal learning; MMRI: Multi-modal representational interaction. All experiments are based on our DeepInteraction framework.
@@
Design choices in the representational interaction encoder. The first row of Table V presents the result of the model without encoder, i.e., two modality-specific representations extracted independently from different backbones are directly fed into the decoder. There is a huge performance gap between this setting and others. This demonstrates that the fusion between heterogeneous modalities is necessary for high-performance 3D detection. To investigate exactly where these improvements come from, we ablate the multi-modal representational interaction (MMRI) and intra-modal representational learning (IML) in the encoder with a different number of layers. We have a couple of observations from Table V: both MMRI and IML contribute to the performance, while the MMRI is more essential. Besides, stacking the encoder layers is essentially better than the shallow interaction.
@@
Figure 5: Illustrations of the heatmaps predicted from BEV representations before (top) and after (bottom) representational interactions. All samples are from the nuScenes val split. (a) A case of the occluded tiny object. (b) A case of a small object at a long distance. (c) An example of two adjacent barriers that are connected together in LiDAR point clouds and thus it is difficult to have their instance-level understanding without the help of visual clues.
@@ V-D3 Ablations on the decoder
TABLE VI: Ablation on the decoder design. We compare the performance between different types of operation employed for the interaction in decoding.
@@
TABLE VII: Ablation on the number of queries used for 3D detection. All experiments are conducted on the DeepInteraction framework.
@@
Multi-modal predictive interaction layer vs. standard DETR [22] prediction. In Table VI, we evaluate the effect of the design for predictive interaction by comparing our multi-modal predictive interaction (MMPI) with standard DETR [22] decoder layer. Note the latter setting means the vanilla cross-attention is used to aggregate multi-modal information as in Transfusion [32]. We further test a mixing design: using the cross-attention for aggregating features in LiDAR representation and MMPI for image representation. The best performance comes from deploying our MMPI for both modalities. The performance gain can be boiled down to the MMPI’s ability to adaptively focus on the local regions of interest, as opposed to the naive cross-attention mechanism that attends to global features.
@@
Since our query embeddings are initialized in a non-parametric and input-dependent manner as in  [32], the number of queries is adjustable during inference. In Figure VII, we evaluate different combinations of query numbers used in training and testing on the DeepInteraction. Overall, the performance is stable over different choices with 200/300 for training/test as the best practice.
@@ V-D4 Ablation on LiDAR backbones
TABLE VIII: Comparison for 3D detection with various point cloud backbones.
@@
We examine the generalization ability of our framework with two different LiDAR backbones: PointPillars [25] and VoxelNet [23]. For PointPillars, we set the voxel size to (0.2m, 0.2m) while keeping the remaining settings as default. For a fair comparison, we use the same number of queries as TransFusion [32]. As shown in Table VIII, due to the proposed multi-modal interaction strategy, DeepInteraction exhibits consistent improvements over the LiDAR-only baseline using either backbone (by 5.5% mAP for the voxel-based backbone, and 4.4% mAP for the pillar-based backbone). These results manifest the generalization ability of our DeepInteraction across varying point cloud backbones. Critically, the improved interaction mechanism is particularly effective for the poor representations extracted from light LiDAR backbone, exhibiting stronger utility on the pillar backbone.
@@ V-D5 Performance breakdown of each category
TABLE IX: Comparison with the LiDAR-only baseline Transfusion-L [32] on nuScenes val split. The mAP breakdown over categories is provided to demonstrate the improvement more comprehensively. “C.V.” and “T.C.” are abbreviations for “construction vehicle” and “traffic cone”.
To demonstrate more fine-grained performance analysis, we compare our DeepInteraction frameworks with the LiDAR-only baseline Transfusion [32] at the category level in terms of mAP on nuScenes val set. We can see from Table IX that our fusion approach achieves remarkable improvements in all categories, especially in tiny or rare object categories.
@@ V-D6 Component analysis of DeepInteraction++
Transformer architecture with deformable attention. In Section III-A, we propose to instantiate representational interaction with a pair of parallel Transformers and replace the original stand-alone attention [55] used in intra-modal interaction with deformable attention [54]. Comparing the a)-b) in Figure 8, we can see that this modification effectively enhances both performance and efficiency. We consider that the performance gain may benefit from the more flexible receptive field of deformable attention, while the efficiency improvement is derived from the highly optimized Transformer implementation.
@@
Scaling backbone and training schedule. In Figure 8 d)-e), we report the additional performance improvements brought by scaling the image backbones. It is worth noting that the improved representational interaction in DeepInteraction++ can further unleash the more powerful representation brought by the scaled backbone and achieve greater marginal gains than the original version as shown in Table II. Furthermore, the revised interaction structure mitigates the overfitting effect, allowing us to further push performance by extending the training schedule as shown in Figure 8 f).
TABLE X: Comparison of end-to-end planning performance with state-of-the-art methods on the nuScenes val set.
@@
TABLE XI: Comparison of the perception and prediction results with VAD-base on the nuScenes val set.

### v3

@@ I Introduction
To overcome the aforementioned limitations, in this work a novel modality interaction strategy, termed DeepInteraction++, for integrating information from different sensors is introduced (Figure 1(b)). Our key idea is to learn and maintain multiple modality-specific representations instead of deriving a single fused representation. This approach enables inter-modality interaction, allowing for the spontaneous exchange of information and the retention of modality-specific strengths with minimal interference between them. Specifically, we start by mapping 3D point clouds and 2D multi-view images into the multi-scale LiDAR BEV features and perspective camera features using two separate feature backbones in parallel. Subsequently, with an encoder we interact heterogeneous features for progressive representation learning and integration in a bilateral manner. To fully exploit per-modality representations, we design a decoder to conduct iteratively multi-modal predictive interaction to yield more accurate perception results.
@@ A.2 Multi-modal representational interaction (MMRI)
where $\boldsymbol{h}^{[i,j]}$ denotes indexing the element at location $(i,j)$ on the 2D representation $\boldsymbol{h}$, and $f_{{\phi}_{p\rightarrow c}}\left(\boldsymbol{h}_{c},\boldsymbol{h}_{p}\right)$ is LiDAR-to-image representational interaction (MMRI-I2L), yielding the image features augmented with the LiDAR information.
@@
Furthermore, we employ the flash attention [52, 53] to minimize the additional computation and memory overhead introduced by this module. The experimental results in Section V demonstrate that this operation provides a beneficial dense context, which complements the original object-centric sparse interaction, thus significantly enhancing detection performance and enabling the extension to end-to-end planning.
@@ A.3 Intra-modal representational learning (IML)
Beyond directly incorporating information from heterogeneous modalities, intra-modal reasoning is helpful for more comprehensive integration of these representations. Therefore, in each layer of the encoder, we conduct intra-modal representational learning complementary to multi-modal interaction. In this work, we utilize deformable attention [54] for intra-modal representational learning, replacing the stand-alone attention [55] in the original version. Considering the scale variance introduced by perspective projection, interaction operation with a more flexible receptive field would be more reasonable than conducting cross-attention within fixed local neighbors as [10]. This modification maintains the original efficient local computation while achieving a more flexible receptive field and facilitating the multi-scale interaction.
@@ A.4 Efficient interaction with grouped sparse attention
To tackle this issue, we carefully examine the distribution of the number of valid image tokens per pillar and divide these pillars into several intervals =$\mathcal{I}=\left\{(N_{i},N_{i+1})\right\}_{i=0}^{N_{\mathrm{interval}}}$. Then we batchify pillars within each interval by padding the number of keys and values to the interval’s upper limit $N_{i+1}$ for attention computations. With careful selection of interval boundaries, this modification significantly reduces memory consumption with negligible impact on parallelism. Besides, it is computationally equivalent to the original implementation, as the padded tokens are masked during the attention process.
@@ III-B Decoder: Multi-modal predictive interaction
TABLE I: Comparison with state-of-the-art methods for 3D object detection on the nuScenes test set. Metrics: mAP(%), NDS(%). † denotes test-time augmentation is used.
@@ IV DeepInteraction++ for end2end autonomous driving
To further demonstrate the scalability and superiority, we extend our DeepInteraction++ to an end-to-end multi-task framework, simultaneously resolving scene perception, motion prediction, and ego-planning tasks. Instead of involving numerous sub-tasks of the driving scenario, we affiliate three additional downstream tasks (including map segmentation, prediction, and planning) following VAD [47], a relatively lightweight framework. Hence, our end-to-end variant can effectively alleviate the memory overhead caused by the complicated interaction encoder and further unleash the multi-task capabilities benefiting from multi-modal representations.
@@ V-A Experimental setup
TABLE II: Run time comparison measured on an NVIDIA RTX A6000 GPU. If not specified with , the performance is evaluated on the nuScenes val set.
@@
TABLE III: Quantitative comparison of detection performance between DeepInteraction and DeepInteraction++ under different image backbones. The results are evaluated on the nuScenes val split.
[…]
backbone
DeepInteraction R50 72.6 69.9 26.7 25.0 27.6 24.8 18.9
DeepInteraction++ 72.9 70.1 26.5 25.1 26.6 24.5 18.9
DeepInteraction Swin-Tiny 72.6 70.0 27.0 25.2 28.1 24.5 18.9
DeepInteraction++ 73.3 70.6 26.8 25.3 26.2 23.4 18.6
@@ V-C Comparison to the state of the arts
Main results. We compare with state-of-the-art alternatives on both the val and test splits of nuScenes dataset. As shown in Table I, our vanilla DeepInteraction has surpassed all its prior arts under the same settings by a considerable margin, and our DeepInteraction++ achieves new state-of-the-art performance with the improved architectural design. Notably, compared to Transfusion [32], which is a representative uniliteral fusion baseline, our DeepInteraction provides a significant performance gain of 2.4% mAP and 1.3% NDS using the same modality-specific backbone and training recipe, verifying the advantages of our multi-modal interaction approach. We provide the per-category results in Table XI. The qualitative results are shown in Figure 4.
Our DeepInteraction++ by default employs a stronger image backbone. To demonstrate that the improvements brought by the revised architecture are consistent and essential, we additionally provide a systematic and comprehensive comparison between DeepInteraction and DeepInteraction++ under the same image backbone on the nuScenes val set. The results in Table III suggest that DeepInteraction++ with a more meticulously designed architecture consistently beats the baseline across most metrics under all settings while adhering to the same hierarchical modality interaction build.
@@
Runtime. We compare the inference speed of all methods on NVIDIA RTX A6000 GPU. As shown in Table II, our method achieves the best performance with faster inference speed than alternative painting-based [2] and query-based [38, 59] fusion approaches. This demonstrates that our method achieves a better trade-off between performance and efficiency. Specifically, feature extraction for multi-view high-resolution camera images contributes the most of the overall latency in a multi-modal 3D detector as verified in [2]. Our interaction modules are built with a relatively lighter model architecture that offers better running speed. From Figure 6, we observe that increasing the number of decoder layers only brings negligible extra latency, which concurs with the same conclusion.
@@ V-D1 Ablations of the modality interaction
Effects of the representational interaction. To demonstrate the superiority of our multi-modal representational interaction, we compare it with a degraded baseline, which does not iteratively refine the image features during the representational interactions. For a fair comparison, both methods use the same number of encoder layers as well as the same decoder. As shown in Table IV a) and c), our representational interaction is more effective than the unilateral fusion alternatives. Besides, we compare the representative Transfusion [32] with the conventional modality fusion strategy in Tables XII, indicating the advantages of our bilateral modality interaction strategy.
@@
Impact of the form of image representation.
TABLE V: Ablation on the image representation. All experiments are based on our DeepInteraction++ framework.
[…]
a) Fused 69.2 72.1
b) BEV 69.4 72.2
c) Perspective 70.3 73.0
During the interactions in both the encoder and decoder, we consistently utilize perspective-form image features instead of converting them to the 3D space beforehand. This design choice is based on two key considerations: (i) With the assistance of LiDAR point clouds, perspective image representation can already achieve precise interaction with LiDAR BEV features, limiting the potential benefits of lifting them into 3D space before the fusion encoder. As indicated in [59, 15], maintaining perspective image features is sufficient for 3D object detection task. (ii) Due to the sparsity of LiDAR data and potential misalignment, this transformation process may be inaccurate and bring irreversible information loss.
To validate this, the comparison of different image representations is demonstrated in Table V. The BEV-form image representation is scattered by the perspective image feature using $\mathcal{M}_{c\rightarrow p}$. In a), the BEV image representation is concatenated with the LiDAR representation and fed into a single-stream Transformer encoder with only IML. In b), the BEV image representation is still kept separate from the LiDAR representation, while the cross-modal interaction in the encoder is replaced by simple deformable attention between two spatially aligned representations. The polar attention is disabled in both b) and c) for a fair comparison. It can be observed that although b) is slightly better than directly fusing them, it still lags significantly behind c) where the image features are kept in perspective. These observations corroborate the rationale behind our design choice.
@@ V-D2 Ablations on the encoder
TABLE VI: Ablation on the encoder design. IML: Intra-modal learning; MMRI: Multi-modal representational interaction. All experiments are based on our DeepInteraction++ framework.
@@
70.6 73.3
TABLE VII: Ablation on the Polar Attention on nuScenes and Waymo dataset based on DeepInteraction++.
nuScenes Waymo (1/5 train)
[…]
w/o polar 70.32 73.02 66.95 61.68
w/ polar 70.63 73.27 67.02 61.70
Design choices in the representational interaction encoder. The first row of Table VI presents the result of the model without encoder, i.e., two modality-specific representations extracted independently from different backbones are directly fed into the decoder. Although this setting has already surpassed LiDAR-only baseline by a considerable margin, there is still a huge performance gap between it and configurations in other rows, underscoring the necessity of representational fusion between heterogeneous modalities for high-performance 3D detection. To investigate exactly where these improvements come from, we ablate the multi-modal representational interaction (MMRI), intra-modal representational learning (IML), and LiDAR-guided cross-plane polar attention (Polar Attn.) in the encoder with various numbers of layers.
We can draw several observations from Table VI: (i) All three components contribute to the performance, while the MMRI and Polar Attn. play more critical roles since they introduce essential inter-modal information exchange. (ii) Stacking more encoder layers is essentially better than the shallow interaction. (iii) While the polar ray attention brings considerable improvement to MMRI, it is insufficient to replace the role of MMRI when used independently. A plausible reason is that although it offers beneficial global context to the original object-centric sparse interaction, it is difficult to provide the precise interaction on its own, which is crucial for 3D object detection. We also evaluate this component on Waymo Open Dataset [71], as shown in Table VII. It demonstrates that indeed, polar attention is more effective for sparse LiDAR datasets than more dense cases (e.g., Waymo Open Dataset).
@@
Figure 5: Illustrations of the heatmaps predicted from BEV representations before (top) and after (bottom) representational interactions. All samples are from the nuScenes val split. (a) Occluded tiny objects. (b) Small objects at long distance. (c) Adjacent barriers connecting together in LiDAR point clouds thus difficult to discriminate without the help of visual clues.
@@ V-D3 Ablations on the decoder
TABLE VIII: Ablation on the decoder design. We compare the performance between different types of operation employed for the interaction in decoding.
@@
TABLE IX: Ablation on the number of queries used for 3D detection. All experiments are conducted on the DeepInteraction framework.
@@
Multi-modal predictive interaction layer vs. standard DETR [22] prediction. In Table VIII, we evaluate the effect of the design for predictive interaction by comparing our multi-modal predictive interaction (MMPI) with standard DETR [22] decoder layer. Note the latter setting means the vanilla cross-attention is used to aggregate multi-modal information as in Transfusion [32]. We further test a mixing design: using the cross-attention for aggregating features in LiDAR representation and MMPI for image representation. Note that this ablation only verifies the module advantage of MMPI layers over DETR layers, where the interaction order with each modality is consistent in all three settings. The best performance comes from deploying our MMPI for both modalities. The performance gain can be boiled down to the MMPI’s ability to adaptively focus on the local regions of interest, as opposed to the naive cross-attention mechanism that attends to global features.
TABLE X: Ablation on alternate interaction in the decoder. The experiments are conducted on the DeepInteraction++ framework.
[…]
LiDAR-then-BEV 72.8 70.4 27.0 25.3 29.3 23.7 18.7
Alternate 73.3 70.6 26.8 25.3 26.2 23.4 18.6
TABLE XI: Comparison with the LiDAR-only baseline Transfusion-L [32] on nuScenes val split. The mAP breakdown over categories is provided to demonstrate the improvement more comprehensively. “C.V.” and “T.C.” are abbreviations for “construction vehicle” and “traffic cone”.
Alternate interaction. The decoder introduced in Section 3 alternately aggregates features from two modalities to maximize their utilization. To validate the effectiveness of this design, we compare it with a non-alternate design where the first three layers access only LiDAR features, followed by image features in the subsequent layers. The results in Table X show that our alternate interaction design yields a considerable advantage on mAOE and NDS, demonstrating its superiority in effectively aggregating multi-modal representation for decoding object attributes.
@@
Since our query embeddings are initialized in a non-parametric and input-dependent manner as in  [32], the number of queries is adjustable during inference. In Figure IX, we evaluate different combinations of query numbers used in training and testing on the DeepInteraction. Overall, the performance is stable over different choices with 200/300 for train/test as the best practice.
@@ V-D4 Ablation on LiDAR backbones
TABLE XII: Comparison for 3D detection with various point cloud backbones.
@@
We examine the generalization ability of our framework with two different LiDAR backbones: PointPillars [25] and VoxelNet [23]. For PointPillars, we set the voxel size to (0.2m, 0.2m) while keeping the remaining settings as default. For a fair comparison, we use the same number of queries as TransFusion [32]. As shown in Table XII, due to the proposed multi-modal interaction strategy, DeepInteraction exhibits consistent improvements over the LiDAR-only baseline using either backbone (by 5.5% mAP for the voxel-based backbone, and 4.4% mAP for the pillar-based backbone). These results manifest the generalization ability of our DeepInteraction across varying point cloud backbones. Critically, the improved interaction mechanism is particularly effective for the poor features extracted from light LiDAR backbone, exhibiting a stronger effect on the pillar backbone.
@@ V-D5 Performance breakdown of each category
To demonstrate more fine-grained performance analysis, we compare our DeepInteraction frameworks with the LiDAR-only baseline Transfusion [32] at the category level in terms of mAP on nuScenes val set. We can see from Table XI that our fusion approach achieves remarkable improvements in all categories, especially in tiny or rare object categories.
@@ V-D6 Component analysis of DeepInteraction++
Transformer architecture with deformable attention. In Section III-A, we propose to instantiate representational interaction with a pair of parallel Transformers and replace the original stand-alone attention [55] used in IML with deformable attention [54]. Comparing the a)-b) in Figure 7, we can see that this modification effectively enhances both performance and efficiency. We consider that the performance gain may benefit from the more flexible receptive field of deformable attention, while the efficiency improvement is derived from the highly optimized Transformer implementation.
@@
Scaling backbone and training schedule. In Figure 7 d)-e), we report the additional performance improvements brought by scaling the image backbones. It is worth noting that the improved representational interaction in DeepInteraction++ can further unleash the more powerful representation brought by the scaled backbone and achieve greater marginal gains than the original version as shown in Table III. Furthermore, the revised interaction structure mitigates the overfitting effect, allowing us to further push performance by extending the training schedule as shown in Figure 7 f).
Figure 9: Failure cases of detection (top) and planning (bottom). The missed ground truth box is highlighted in red.
TABLE XIII: Comparison of end-to-end planning performance with state-of-the-art methods on the nuScenes val set.
@@
TABLE XIV: Comparison of the perception and prediction results with VAD-base on the nuScenes val set.